A 50 mL graduated cylinder contains 25.0 mL of water. A 42.5040 g piece of gold is placed in the graduated cylinder and the water level rises to 27.2 mL. What is the density of the piece of gold? (a) 19.3 g/mL (b) 15.6 g/mL (c) 5.90 g/mL (d) 12.3 g/mL (e) 3.63 g/mL
Answer #1

initial volume of water = 25.0 mL

volume of water after piece of gold added = 27.2 ml

volume of gold peace = 27.2 - 25.0 = 2.2 mL

mass of gold = 42.5040 g

density = mass / volume

density = 42.5040 / 2.2

density = 19.3 g/mL

answer = option a = 19.3 g/mL